OpenStudy (anonymous):

how do you find all the zeros of: f(x)=5x^2+27x+10

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

this one factors as \[(5x+2)(x+5)\] \[(5x+2)(x+5)=0\] \[5x+2=0\] \[x=-\frac{2}{5}\] or \[x+5=0\] \[x=-5\]